OpenStudy (anonymous):

m^3 n^2 p^5 _____________ m^4 np^2

OpenStudy (saifoo.khan):

Alright, so let's solve.

OpenStudy (saifoo.khan):

Use this: \[\Large \frac{a^b}{a^c} \to a^{b-c}\]

OpenStudy (anonymous):

that's what i'm using

OpenStudy (saifoo.khan):

\[\Large \frac{m^3 n^2 p^5 }{m^4 np^2 } \to m^{3-4}n^{2-1}p^{5-3}\]

OpenStudy (anonymous):

m-1 n1 p3 m=1/p3 (the number one can not be raised ????)

OpenStudy (saifoo.khan):

\[\Large m^{-1}n^1 p^2 \to \frac{np^3}{m}\]

OpenStudy (anonymous):

this takes alot of practice ..... my question to you is how is the np3 on top is because we had raise it from the bottom

OpenStudy (saifoo.khan):

We have to get our final answer in +ve exponents.. That's how our fowmula works.